Two vectors, A and B are added together to form the resultant R = 3ˆı − 4 jˆ. If

⃗⃗ vector A = 5ˆı − 2 jˆ what is vector B?
Answer #1

ro Given R = 3î - 4 Ā = 51-2 B = ? .28 As = R B= (RCA) per question à to = S. B - (07-41)-(s{ -23) = -2î - 2 S. B = -21 -23